Please help me solve this question:
A game involves rolling a die once. If a number less than 3 appears,you lose $60. If a 4 or 6 appears,you win $120. If a 5 appears,nothing happens.
A.) show the probability distribution for the money.
B.) calculate the expected monetary value of the game.

Answer by Theo(13342)

your probabilities are incomplete.
there is no condition for a value of 3 being chosen.
i will assume that if a 3 or a 5 appears, nothing happens.
that makes the probabilities come out correctly.
p(<3) = 2/6
p(4 or 6) = 2/6
p(3 or 5) = 2/6
your expected value is:
2/6 * -60 + 2/6 * 120 + 0 which is equal to -120/6 + 240/6 = 120/6 = 20.
on the average, if you play the game an infinite number of times, you should collect 20 dollars each time.
example:
you play the game 1200 times.
2/6 * 1200 = 2400/6 = 400 times you will lose 60 dollars.
400 times you will win 120 dollars
400 times you will get nothing.
your total winning is 400 * 120 = 48,000.
your total losing is 400 * 60 = 24,000
your net gain is 24,000
divide that by 1200 and you have average a net gain of 20 per game.